OpenStudy (anonymous):

Calculate the value of x side AC is the diameter. A. 9 B. 18 C. 17 D. 152

OpenStudy (anonymous):



OpenStudy (johnweldon1993):

well the degrees of a circle add up to 360 right? this is half...so 180 this means your 3 degrees...add up to 180 6x + 5 + 5x + 3 + 8x + 1 = 180 combine like terms 6x + 5x + 8x = 19x 5 + 3 + 1 = 9 19x + 9 = 180 -9 -9 19x = 171 ---------- 19 19 x = 9